# Bulk condensation by an active interface

### v3

@@ Acknowledgements.
We have benefited from valuable discussions with M Barma, M E Cates, A Maitra, R Mondal, and M Rao. SR thanks the SERB, India for a J C Bose Fellowship, ICTS-TIFR for a Simons Visiting Professorship, the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Anti-diffusive dynamics: from sub-cellular to astrophysical scales where work on this paper was undertaken, supported by EPSRC grant no EP/R014604/1, and ICTS-TIFR for support and hospitality during Active Matter and Beyond. RK was supported by the UGC, India, AKS acknowledges a National Science Chair Professorship of the DST, Government of India, and HS thanks the SERB, India, for support under grant no. SRG/2022/000061-G.

### v4

@@ Acknowledgements.
We have benefited from valuable discussions with M Barma, M E Cates, A Maitra, R Mondal, C Nardini, and M Rao. SR thanks the SERB, India for a J C Bose Fellowship, ICTS-TIFR for a Simons Visiting Professorship, the Isaac Newton Institute for Mathematical Sciences, Cambridge, for a Rothschild Distinguished Visiting Fellowship and support and hospitality during the programme Anti-diffusive dynamics: from sub-cellular to astrophysical scales where work on this paper was undertaken, supported by EPSRC grant no EP/R014604/1, and ICTS-TIFR for support and hospitality during Active Matter and Beyond. RK was supported by the UGC, India, AKS acknowledges a National Science Chair Professorship of the DST, Government of India, and HS thanks the SERB, India, for support under grant no. SRG/2022/000061-G.